OpenStudy (anonymous):

WILL MEDAL AND FAN ALL WHOS RIGHT 1/5(2x - 15) + 3/5x = 4 1/4

OpenStudy (anonymous):

4

OpenStudy (deadpoolroxz):

yeah 4

OpenStudy (anonymous):

just 4 ? its suppose to be 1 whole number then a fraction

OpenStudy (anonymous):

x in (-oo:+oo) (1/5)*(2*x-15)+(3/5)*x = 1-4+4 // - 1-4+4 (1/5)*(2*x-15)+(3/5)*x-4-1+4 = 0 1/5*(2*x-15)+(3*x)/5-4-1+4 = 0 (1/5*5*(2*x-15))/5+(3*x)/5+(-4*5)/5+(-1*5)/5+(4*5)/5 = 0 1/5*5*(2*x-15)+3*x-4*5-1*5+4*5 = 0 5*x-20-15-5+20 = 0 5*x-35-5+20 = 0 5*x-40+20 = 0 5*x-20 = 0 (5*x-20)/5 = 0 (5*x-20)/5 = 0 // * 5 5*x-20 = 0 5*x-20 = 0 // + 20 5*x = 20 // : 5 x = 20/5 x = 4 x = 4

OpenStudy (deadpoolroxz):

Solution for 1/5(2x-15)+3/5x=4+1-4 equation: x in (-oo:+oo) (1/5)*(2*x-15)+(3/5)*x = 1-4+4 // - 1-4+4 (1/5)*(2*x-15)+(3/5)*x-4-1+4 = 0 1/5*(2*x-15)+(3*x)/5-4-1+4 = 0 (1/5*5*(2*x-15))/5+(3*x)/5+(-4*5)/5+(-1*5)/5+(4*5)/5 = 0 1/5*5*(2*x-15)+3*x-4*5-1*5+4*5 = 0 5*x-20-15-5+20 = 0 5*x-35-5+20 = 0 5*x-40+20 = 0 5*x-20 = 0 (5*x-20)/5 = 0 (5*x-20)/5 = 0 // * 5 5*x-20 = 0 5*x-20 = 0 // + 20 5*x = 20 // : 5 x = 20/5 x = 4 x = 4

OpenStudy (anonymous):

that makes no since

OpenStudy (anonymous):

just take the steps 1 by 1 i just got off this lesson so 4 is correct

OpenStudy (anonymous):

1/5(2x - 15) + 3/5x = 4 (1/4) 2/5x - 15/5 + 3/5x = 4(1/ 4) 5/5 x - 3 = 17/4 x = 17/4 -3 x = 17 - 12/4 x = 5/4 x = 1(1/4)

OpenStudy (deadpoolroxz):

the answer is four use calculater

OpenStudy (anonymous):

the answer was 7 1/4 ... and i failed the test thing 3 times already... @deadpoolroxz @sangya21

OpenStudy (deadpoolroxz):

sarry

OpenStudy (anonymous):

Sorry. :(

OpenStudy (deadpoolroxz):

yeah bro i was harh hope ur teacher puts it back on

OpenStudy (anonymous):

1/5(2x - 15) + 3/5x = 4 (1/4) 2/5x - 15/5 + 3/5x = 4(1/ 4) 5/5 x - 3 = 17/4 I made mistake in next step instead of - there should be plus sign x = 17/4 +3 x = 17 + 12/4 x = 29/4 x = 7(1/4)

OpenStudy (anonymous):

@BillNye01 please can you upload a picture of the question?The expression can suggest many expression.s

OpenStudy (anonymous):

@Princer_Jones ty for trying to help but im done with this one